# Hydrodynamic flows induced by localized torques (rotlets) in wedge-shaped geometries

### v1

@@ VI Conclusions
Summarizing, we here derived analytical expressions for low-Reynolds-number fluid flows induced by a localized torque in a wedge-shaped geometry. No-slip surface conditions are applied at the surfaces of the wedge-shaped confinement. To perform the mathematical derivation, we rely on the Fourier-Kontorovich-Lebedev transform. From the illustration of resulting flow fields, we observe the obvious and expected distortion of the resulting fluid flows and vortices by the wedge-shaped geometry.
@@ Acknowledgements.
E.T. acknowledges support from the Engineering and Physical Sciences Research Council (EPSRC) under grant no. EP/W027194/1. J.M. and A.M.M. thank the Deutsche Forschungsgemeinschaft (German Research Foundation, DFG) for support through project no. 541972050 (DFG reference no. ME 3571/12-1).
@@
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.

### v2

@@ VI Conclusions
Summarizing, we here derived analytical expressions for low-Reynolds-number fluid flows induced by a localized torque in a wedge-shaped geometry. No-slip surface conditions are applied at the surfaces of the wedge-shaped confinement. To perform the mathematical derivation, we rely on the Fourier-Kontorovich-Lebedev transform. From the illustration of resulting flow fields, we observe the expected distortion of the resulting fluid flows and vortices by the wedge-shaped geometry.
@@
The point-particle approximation employed in the present work is generally valid in the far-field limit, when the particle is sufficiently distant from the nearest interface. A systematic assessment of its accuracy would require either exact analytical solutions for fully extended particles or detailed numerical simulations. Nevertheless, based on our previous experience with planar and spherical interfaces, our approximation—despite its simplicity—often provides remarkably accurate predictions of hydrodynamic mobilities, even at distances comparable to the particle diameter, which highlights its usefulness for practical calculations, qualitative estimates at smaller distances, and theoretical analyses [75, 76, 77]. We note that our theory may exhibit quantitative deviations before near-field lubrication forces become significant. Nevertheless, when the particle approaches one of the two flat surfaces of the wedge, at sufficient distance from the tip of the wedge and the other surface, lubrication interactions with this one surface dominate. In this case, the approximation valid for one flat surface provides a reasonable leading-order description. In future work, it would be interesting to study the problem using exact analytical frameworks, such as bipolar coordinates or related methods.
@@ Acknowledgements.
J.M. and A.M.M. thank the Deutsche Forschungsgemeinschaft (German Research Foundation, DFG) for support through project no. 541972050 (DFG reference no. ME 3571/12-1). E.T. acknowledges support from the Engineering and Physical Sciences Research Council (EPSRC) under grant no. EP/W027194/1.
@@
Mathematica scripts that compute the flow field at a specified coordinate are available on the Zenodo data repository at https://zenodo.org/records/18350815. All other data are available from the corresponding author upon reasonable request.